# A Survey on Optimization Studies of Group Centrality Metrics

### v1

@@ 3 Brief history on group centrality
In 2022, Camur et al. (2022) proposes a more effective formulation along with an exact solution technique (i.e., Benders Decomposition (Rahmaniani et al., 2017)) for the SDC problem, proposed by Vogiatzis and Camur (2019). They also present types of networks where the SDC problem is solvable in polynomial time. Additionally, in Rasti and Vogiatzis (2022) the authors investigate stars, cliques, and representative structures with respect to the same three group centrality metrics. The authors further define a representative set as a group of nodes centered at a node; in other words, it can be viewed as a relaxation of the star, allowing one node to serve as a center while the leaves may or may not share edges between each other. They also propose CBB algorithms for all three structures and all three metrics. Furthermore, Rajbhandari et al. (2022) revisits the group closeness maximization problem for directed graphs that do not require connectedness, proposing a heuristic algorithm.
@@
Unsurprisingly, in 2023, we witness even more studies in group centrality research. To start with, Camur et al. (2023) integrates probability values regarding edge connectivity and introduce the very first “probabilistic” group centrality metric as stochastic pseudo-star degree centrality. According to the authors a pseudo-star is a set of nodes forming a star with at least a probability $\theta\in\left[0,1\right]$. This probability can be viewed as the equivalent of the probability of all edges between the designated center and stars to exist, as well as no edges between any two leaves existing. We note here the similarity between the pseudo-star and a probabilistic clique (Miao et al., 2014). A probabilistic clique is defined as a set of nodes that form a clique with at least probability $\theta\in\left[0,1\right]$.
@@
Returning to the pseudo-star degree centrality problem (Camur et al., 2023), this novel group centrality metric is shown to be more effective than its deterministic correspondent (i.e., SDC, see Vogiatzis and Camur (2019)) in determining essential proteins in protein-protein interaction networks. That said, it is important to note that this study does not employ any stochastic optimization modeling or techniques. On the contrary, the authors utilize a priori available probability values (obtained in the context of the practical application from protein interaction frequencies) to identify the highest weighted degree pseudo-star.
@@
Then, Matsypura et al. (2023) proposes other novel group centrality metrics called most degree-central walks and paths. The study indicates that while the most central shortest paths can be identified in polynomial time, most central walks (e.g., paths) with a pre-defined length remains $\mathcal{NP}$-complete. It is important to notice that representative sets of nodes, as introduced in Rasti and Vogiatzis (2022), do not cover walks and paths. The authors impose a specific connectivity constraints among the nodes in the subgraph. Thus, this can be classified as a distinct motif different from a clique and a star (see Figure 2(a) and (b), compared to (c) and (d), respectively).
@@
After Chen et al. (2016) introduces group closeness centrality from an optimization perspective, Staus et al. (2023) propose two exact solution techniques for the same problem. Lastly, Lagos et al. (2023) revisits the problem of identifying groups with maximum betweenness centrality as previously discussed by Puzis et al. (2007) and Veremyev et al. (2017). The authors integrate mathematical modeling with randomized path sampling.
@@
Before moving into the next section where we continue our discussion with analyzing group structures and optimization methods utilized, we would like to mention one more aspect of group centrality. During our review, we encountered a group of work in which group centrality is addressed using the Shapley value (Hart, 1989). This set of work is predominantly considered in social network analysis using game theory, which is another branch of OR (see (Szczepanski et al., 2012; Michalak et al., 2013; Flores et al., 2016; Sun et al., 2020; Belik, 2023) in order). However, these research products seem to be more independent of the framework that we have presented above. While all of these studies reference Everett and Borgatti (1999) regarding group centrality, we did not find any reference to the rest of the works presented in Figure 4. Thus, we prefer not to include these studies during our analysis in the next section.
@@ 4 Structures and computational approaches
Group
Chen et al. (2016); Ohara et al. (2017); Zhao et al. (2017); Bergamini et al. (2018); Veremyev et al. (2019); Li et al. (2019); Angriman et al. (2019, 2021); Rajbhandari et al. (2022); Staus et al. (2023)
Puzis et al. (2007); Brandes (2008); Dolev et al. (2009, 2010); Fink and Spoerhase (2011); Mahmoody et al. (2016); Veremyev et al. (2017); Akgün and Tural (2020); Lagos et al. (2023)
Borgatti (2006); Fushimi et al. (2019); Angriman et al. (2020); Li et al. (2020)
Clique
Vogiatzis et al. (2015); Pastukhov et al. (2018); Zhong et al. (2021)
Vogiatzis et al. (2015); Nasirian et al. (2020)
Vogiatzis et al. (2015); Veremyev et al. (2017); Rysz et al. (2018)
Star
Vogiatzis and Camur (2019); Camur et al. (2022, 2023)
Rasti and Vogiatzis (2022)
Rasti and Vogiatzis (2022)
@@
Rasti and Vogiatzis (2022)
Rasti and Vogiatzis (2022)
Rasti and Vogiatzis (2022)
Walk
Matsypura et al. (2023)
As previously mentioned, groups with the largest degree centrality are more closely related to the maximum coverage problem literature; thus, the corresponding section in Table 1 does not include any of these studies. Extensive research is conducted on groups of nodes without any restrictions in terms of both closeness and betweenness centrality (see the first row in Table 1). However, from OR scientists, we observe significant interest in cohesive groups. Specifically, they focus on particular structures, with cliques attracting the most attention. In recent years, the star structure is also examined from a group centrality perspective. In fact, Rasti and Vogiatzis (2022) generalizes all clique and star-type graph structures under representative sets, covering all the major group centrality metrics in that study. Additionally, the authors define group centrality on a nodal level as the value of the group (inducing a specific structure) with maximum centrality value, so long as it includes that node as a member (in the case of cliques) or as a center (in the case of stars).
Since much of the existing literature covers mostly both star and clique structures, Matsypura et al. (2023) examines walk and path centrality, expanding the concept of cohesive groups. This study opens a new window into the group centrality literature from a special structure perspective. Moreover, we note a study by van der Grinten et al. (2021), which introduces a novel centrality measure named Forest Closeness Centrality for a group of nodes. This concept builds upon the Forest Distance Closeness initially proposed by Jin et al. (2019). However, due to the absence of a formal journal publication on this new measure and its still-developing stage of research, we avoid providing more details about it in our study. Yet, we mention it in the hopes that it will spark an interest in further investigation in very novel structures and centrality metrics.
@@
Vogiatzis et al. (2015); Pastukhov et al. (2018); Vogiatzis and Camur (2019); Zhong et al. (2021); Camur et al. (2022); Rasti and Vogiatzis (2022); Matsypura et al. (2023); Camur et al. (2023)
Vogiatzis et al. (2015); Bergamini et al. (2018); Veremyev et al. (2019); Rahim Taleqani et al. (2020); Nasirian et al. (2020); Rasti and Vogiatzis (2022); Staus et al. (2023)
Vogiatzis et al. (2015); Veremyev et al. (2017); Rysz et al. (2018); Rasti and Vogiatzis (2022)
@@
Fushimi et al. (2019); Matsypura et al. (2023)
Chen et al. (2016); Veremyev et al. (2019); Rahim Taleqani et al. (2020); Rajbhandari et al. (2022)
Puzis et al. (2007); Dolev et al. (2010); Akgün and Tural (2020)
Borgatti (2006); Fushimi et al. (2019)
@@
Vogiatzis and Camur (2019)
Chen et al. (2016); Ohara et al. (2017); Zhao et al. (2017); Bergamini et al. (2018); Li et al. (2019); Angriman et al. (2019, 2021)
Dolev et al. (2009); Fink and Spoerhase (2011); Mahmoody et al. (2016); Rysz et al. (2018)
Angriman et al. (2020)
@@
Vogiatzis et al. (2015); Pastukhov et al. (2018); Zhong et al. (2021); Rasti and Vogiatzis (2022)
Nasirian et al. (2020); Rasti and Vogiatzis (2022); Staus et al. (2023)
Rasti and Vogiatzis (2022)
@@
Pastukhov et al. (2018); Camur et al. (2022), Camur et al. (2023)
Next, we examine the optimization methodologies adopted by OR researchers. To begin, we categorize solution methods into five main groups: (a) optimization modeling (e.g., integer programming, linear programming), (b) heuristic algorithms, (c) heuristics with approximation guarantees (i.e., approximation algorithms), (d) combinatorial branch-and-bound (CBB) algorithms (this also includes “smart selection” algorithms that use lower and upper bounds), and (e) decomposition algorithms (e.g., Benders Decomposition). For group (a), it is important to note that studies offering a heuristic approach or an exact solution technique are highly likely to include a mathematical model, leading to potential duplication among cells. Regarding centrality, we report the same metrics as shown in Table 1. As a result, we present Table 2 to categorize recent works based on the methodology adopted.
@@
Conversely, research related to highest degree star centrality frequently employ decomposition algorithms to solve the large-scale models. In both Camur et al. (2022) and Camur et al. (2023), the authors implement various Benders decomposition strategies, including traditional Benders approach, modern Benders, two-phase, three-phase decomposition, and Logical Benders. The problems of star centrality with the highest closeness and betweenness, on the other hand, are tackled by CBB approaches.
It is evident that studies lacking any specific motif are often tackled using heuristic or approximation algorithms. In essence, there are additional studies regarding groups without motifs that we do not include in Figure 4. For instance, the concept of group closeness, studied by Chen et al. (2016), is later adapted by other researchers to improve the approximation guarantee. Bergamini et al. (2018) then scales up the existing approximation, and Angriman et al. (2019) offers an algorithm that is one to two orders of magnitude faster, based on a new local search heuristic. Subsequently, Angriman et al. (2021) provides a constant approximation for the same local search algorithm.
@@
Nevertheless, there are few exceptions where group centrality metrics are applied to real-world scenarios. Specifically, we observe significant applications in biological networks, particularly in protein-protein interaction networks (refer to Vogiatzis and Camur (2019); Camur et al. (2022); Rasti and Vogiatzis (2022); Camur et al. (2023)). These studies utilize group structures such as star and representative sets in analyzing protein-protein interaction networks to determine essential proteins for the survival of the organisms.
@@ 5 Future work
We observe that the OR community has seen significant advancements on group centrality in the last decade as discussed in Section 3. In addition, we provide additional analysis to provide more depth insights in Section 4. Yet, there are several unexplored areas that might of interest to optimization experts. As recently introduced by Matsypura et al. (2023), both walk and path group centrality are evaluated primarily from the largest degree perspective. Thus, there is a research gap for these novel group centrality metrics with respect to closeness and betweenness centrality that we believe needs to be addressed.
Another research direction lies in the domain of tree structures within graphs. Trees, defined as connected graphs where there exists a unique path between each node pair, as a special type of graph, could serve as a basis for developing new group centrality measures. While the concept of spanning tree centrality exists in literature, it differs from our understanding in this paper. The existing studies focus on the probability of an edge being used in a uniformly sampled spanning tree (Teixeira et al., 2013; Hayashi et al., 2016). Identifying “central” trees could serve in application as developing communication protocols; their relationships to connected dominating sets (Guha and Khuller, 1998; Validi and Buchanan, 2020) could also be further investigated.
@@
Newman (2005) introduces a new measure of betweenness centrality based on random walks. This measure considers the contribution of all paths, not just the shortest ones, while placing more weight on the shortest paths and considering generating random walks (see Noh and Rieger (2002)). Later, Gillani et al. (2021) extends this idea to group to group version where centrality of a node is measured through a set of source and a set of sink nodes. They indicate that “We look forward to extend random walk based centrality definition to the group centrality setting in which we measure the centrality of groups rather than individual nodes.”. We believe that this might be another avenue for further research. Even though this metric might sounds similar to the GED-centrality proposed by Angriman et al. (2020), we note that random-walk centrality is more related to algebraic-based metrics including the Laplacian equations.
As for the solution techniques, we note that there is no study proposing exact solution approaches for groups without motifs. Investigating this could lead to significant contributions. Additionally, there is a clear path towards exploring probabilistic group centrality metrics, such as examining the problem of maximum probabilistic clique (Miao et al., 2014) from a group centrality perspective. Assuming static, unchanging conditions in large-scale network can be a significant limitation in real-world scenarios where uncertainty and variability are inherent. That is why exploring this would be important to understand network dynamics better.
Considering that current models in OR literature are predominantly deterministic, introducing stochastic optimization techniques, like two-stage stochastic models, could be a valuable addition to the field. Such models can help in identifying influential groups that are crucial for the robustness and efficiency of a network under probabilistic conditions. This approach could lead to more resilient and adaptive network designs, thereby enhancing decision-making processes in areas such as infrastructure planning, supply chain and logistics. Additionally, considering the fact that for many networks we do not have a full picture of all the existing nodes and edges (see, e.g., illicit networks which have attracted significant recent interest as in the works by Kosmas et al. (2023); Anzoom et al. (2023)), stochastic extensions to group centrality can lead to more realistic identification of key actors in networks with hidden elements.
@@
We start, again, from a very important work by Matsypura et al. (2023): identifying paths and walks of high degree can have applications in using autonomous vehicles to monitor a number of targets (by approaching them). As autonomous vehicles are becoming more and more prominent in applications such as humanitarian relief operations, last mile deliveries, health visits, etc., we believe that the study of closeness metrics from the perspective of walks, paths, and tours/cycles will be an important piece of the puzzle.
Nodal centrality has been prominent in interesting new societal problems, such as investigating criminal and illicit networks (Bichler et al., 2017; Grassi et al., 2019). That said, group centrality metrics present a unique opportunity to identify many key illicit actors at once. Moreover, we mentioned earlier the need for more stochastic measures of group centrality. These would be particularly interesting in this context, due to the large number of hidden or obscured entities and relationships.
Additionally, a number of researchers have investigated criticality and centrality from the perspective of disaster management (mostly evacuation). Continuing on that path, we believe it is important to include refueling constraints (see, e.g., Vogiatzis and Kontou (2024)) in the calculation of group centrality when applied to evacuation, especially considering the scarcity of gas (and soon other alternative fuels) during an emergency. Finally, another avenue for research is in pedagogy and the formation of groups in the classroom (see, for example, Hill and Peuker (2023)). Utilizing group centrality to identify groups with “natural leaders” or groups that satisfy a specific connectivity constraint (e.g., high degrees to other groups or mentors, high closeness to some mentor, etc.) can improve team formation and lead to improved outcomes inside and outside the classroom.

### v3

@@ 3.5 Recent Developments in Group Centrality
Recent years have witnessed significant advancements in the field of group centrality, characterized by the introduction of innovative metrics and refined optimization methods. In 2022, Camur et al. [19] propose a more effective formulation along with an exact solution technique (i.e., Benders Decomposition [79]) for the SDC problem, proposed by Vogiatzis and Camur [99]. They also present types of networks where the SDC problem is solvable in polynomial time. Additionally, Rasti and Vogiatzis [81] investigate stars, cliques, and representative structures with respect to the same three group centrality metrics. The authors further define a representative set as a group of nodes centered at a node; i.e., it can be viewed as a relaxation of the star, allowing one node to serve as a center while the leaves may or may not share edges between each other. They propose CBB algorithms for all three structures and all three metrics. Furthermore, Rajbhandari et al. [80] revisit the group closeness maximization problem for directed graphs that do not require connectedness, proposing a heuristic algorithm.
@@
Unsurprisingly, in 2023, we witness even more studies in group centrality research. To start with, Camur et al. [20] integrate probability values regarding edge connectivity and introduce the very first “probabilistic" group centrality metric as stochastic pseudo-star degree centrality. According to the authors a pseudo-star is a set of nodes forming a star with at least a probability $\theta\in\left[0,1\right]$. This probability can be viewed as the equivalent of the probability of all edges between the designated center and stars to exist, as well as no edges between any two leaves existing. We note here the similarity between the pseudo-star and a probabilistic clique [63]. A probabilistic clique is defined as a set of nodes that form a clique with at least probability $\theta\in\left[0,1\right]$.
@@
Returning to the pseudo-star degree centrality problem [20], this novel group centrality metric is shown to be more effective than its deterministic correspondent (i.e., SDC, see [99]) in determining essential proteins in protein-protein interaction networks. That said, this study does not employ any stochastic optimization modeling or techniques. On the contrary, the authors utilize a priori available probability values (obtained in the context of the practical application from protein interaction frequencies) to identify the highest weighted degree pseudo-star.
@@
Then, Matsypura et al. [60] propose other novel group centrality metrics called most degree-central walks and paths. The study indicates that while the most central shortest paths can be identified in polynomial time, most central walks (e.g., paths) with a pre-defined length remains $\mathcal{NP}$-complete. It is important to notice that representative sets of nodes, as introduced in Rasti and Vogiatzis [81], do not cover walks and paths. The authors impose a specific connectivity constraints among the nodes in the subgraph. Thus, this can be classified as a distinct motif different from a clique and a star (see Figure 2(a) and (b), compared to (c) and (d), respectively).
@@
After Chen et al. [22] introduce group closeness centrality from an optimization perspective, Staus et al. [89] propos two exact solution techniques for the same problem. Lastly, Lagos et al. [54] revisit the problem of identifying groups with maximum betweenness centrality as previously discussed by Puzis et al. [77] and Veremyev et al. [97]. The authors integrate mathematical modeling with randomized path sampling.
3.6 Excluded Aspects of Group Centrality
Throughout the previous subsections, we have mentioned a series of works that are relevant and related, but were excluded from this survey. These included the detection of critical nodes and sets of graph elements, graph partitioning and community detection alongside “cut” problems, facility location problems, and coverage problems on graphs. We provide a list of indicative works and surveys in the next paragraphs to guide researchers that are interested in these areas and their relationships to group centrality.
Critical nodes and graph elements are very related to our survey and focus, which is also the reason why Figure 4 does include the foundational work “Identifying sets of key players in a social network” by Borgatti in 2006 [13]. However, our focus remains firmly on the realm of centrality (which is usually a proxy for importance or criticality), hence we do not fully discuss a series of works in this area. Interested readers will find the surveys by Walteros and Pardalos in 2012 [104] and by Lalou et al. in 2018 [55] particularly useful.
Another part of the related literature includes graph communities. Communities are often graph structures that are of interest in our survey, e.g., in the form of cliques or quasi-cliques. Earlier in this work, we mentioned how betweenness centrality served for the detection of community structures in networks [71]. Newman is also responsible for one of the earliest comprehensive surveys of identifying communities in networks [69]. A very recent and thorough survey on a variety of tools and applications of community detection is by Fortunato [33]. While centrality is often used as a way to identify nodes/edges to consider while detecting communities (see, e.g., Girvan-Newman method and betweenness [71] or Fortunato method and current flow centrality [34]), the focus of these studies is on community detection rather than centrality, so we exclude them from discussion. Sometimes, especially in problems involving community detection through cuts (see, e.g., normalized cuts [88] and other variants), the focus can be on communities of interest to this survey (e.g., the structural requirements in [103]), but there is no mention of centrality, so these are also excluded from our discussion.
Additionally, location and siting problems have long benefited from the use of centrality metrics. Specifically, Angriman et al. mention that group closeness (which is one of the focal group centrality metrics in this survey) appear to be a particularly useful objective in facility location problems [4, 3]. That said, the literature in facility location is vast (see, e.g., the recent surveys in [65, 25]), so we omit these works from our study unless they focus specifically on group centrality metrics with an application to facility location and siting.
Finally, we discuss the relationship of group degree to maximum coverage problems in networks. For example, the study titled “Group Degree Centrality and Centralization in Networks" by Krnc and Škrekovski [53] examines the set of nodes with the largest degree. However, the focus on largest-degree subgraphs without any connectivity or structural restrictions aligns more with the maximum coverage problem (see Megiddo et al. [62], which is beyond our current scope. Similarly, the concept of group coverage centrality, mostly studied by researchers outside the field of OR and out of the scope the work, are not included in our review [108, 61, 28, 51], despite their importance.
@@ 3.6 Excluded Aspects of Group Centrality
Before moving into the next section where we continue our discussion with analyzing group structures and optimization methods utilized, we would like to mention one more aspect of group centrality. During our review, we encountered a group of work in which group centrality is addressed using the Shapley value [107]. This set of work is predominantly considered in social network analysis using game theory, which is another branch of OR (see [93, 64, 32, 92, 9] in order). However, these research products seem to be more independent of the framework that we have presented above. While all of these studies reference Everett and Borgatti [29] regarding group centrality, we did not find any reference to the rest of the works presented in Figure 4. Thus, we prefer not to include these studies during our analysis in the next section.
@@ 4 Structures and computational approaches
Group [22, 73, 109, 10, 98, 56, 5, 3, 80, 89] [77, 15, 27, 26, 31, 59, 97, 2, 54] [13, 40, 4, 57]
Clique [102, 76, 110] [102, 68] [102, 97, 83]
Star [99, 19, 20] [81] [81]
@@
[81] [81] [81]
Walk [60]
As previously mentioned, groups with the largest degree centrality are more closely related to the maximum coverage problem literature; thus, the corresponding section in Table 1 does not include any of these studies. Extensive research is conducted on groups of nodes without any restrictions in terms of both closeness and betweenness centrality (see the first row in Table 1). However, from OR scientists, we observe significant interest in cohesive groups. Specifically, they focus on particular structures, with cliques attracting the most attention. In recent years, the star structure is also examined from a group centrality perspective. In fact, Rasti and Vogiatzis [81] generalize all clique and star-type graph structures under representative sets, covering all the major group centrality metrics in that study. Additionally, the authors define group centrality on a nodal level as the value of the group (inducing a specific structure) with maximum centrality value, so long as it includes that node as a member (in the case of cliques) or as a center (in the case of stars).
Since much of the existing literature covers mostly both star and clique structures, Matsypura et al. [60] examine walk and path centrality, expanding the concept of cohesive groups. This study opens a new window into the group centrality literature from a special structure perspective. Moreover, we note a study by Grinten et al. [96], which introduces a novel centrality measure named Forest Closeness Centrality for a group of nodes. This concept builds upon the Forest Distance Closeness initially proposed by Jin et al. [49]. However, due to the absence of a formal journal publication on this new measure and its still-developing stage of research, we avoid providing more details about it in our study. Yet, we mention it in the hopes that it will spark an interest in further investigation in very novel structures and centrality metrics.
@@
[102, 76, 99, 110, 19, 81, 60, 20] [102, 10, 98, 78, 68, 81, 89] [102, 97, 83, 81]
@@
[40, 60] [22, 98, 78, 80] [77, 26, 2] [13, 40]
@@
[99] [22, 73, 109, 10, 56, 5, 3] [27, 31, 59] [4]
@@
[102, 76, 110, 81] [68, 81, 89] [83, 81]
@@
[76, 19], [20]
Next, we examine the optimization methodologies adopted by OR researchers. We establish a distinction between exact and heuristic algorithms to reflect their different roles and efficacy in solving complex problems. Exact methods systematically explore the solution space to guarantee the identification of optimal solutions. During our review, we encountered three main groups: a) mathematical modeling techniques such as integer programming and linear programming, b) branch-and-bound approaches, and c) decomposition algorithms.
Conversely, heuristic methods, which we segment into simple heuristics and approximation algorithms in this study, prioritize computational efficiency and practical applicability over guaranteed optimality. Simple heuristics provide feasible solutions quickly without assurances on closeness to optimality. Approximation algorithms, as a specialized form of heuristics, offer a balance by providing solutions with known bounds on their deviation from optimality. This categorization aids in understanding the trade-offs between solution quality and computational resources, thereby guiding the selection of appropriate methods based on specific problem requirements and constraints.
In light of these, we categorize solution methods into five main groups: (a) optimization modeling (e.g., integer programming, linear programming), (b) heuristic algorithms, (c) heuristics with approximation guarantees (i.e., approximation algorithms), (d) combinatorial branch-and-bound (CBB) algorithms (this also includes “smart selection” algorithms that use lower and upper bounds), and (e) decomposition algorithms (e.g., Benders Decomposition). For group (a), it is important to note that studies offering a heuristic approach or an exact solution technique are highly likely to include a mathematical model, leading to potential duplication among cells. Regarding centrality, we report the same metrics as shown in Table 1. As a result, we present Table 2 to categorize recent works based on the methodology adopted.
@@
Conversely, research related to highest degree star centrality frequently employ decomposition algorithms to solve the large-scale models. In both Camur et al. [19] and Camur et al. [20], the authors implement various Benders decomposition strategies, including traditional Benders approach, modern Benders, two-phase, three-phase decomposition, and Logical Benders. The problems of star centrality with the highest closeness and betweenness, on the other hand, are tackled by CBB approaches.
It is evident that studies lacking any specific motif are often tackled using heuristic or approximation algorithms. In essence, there are additional studies regarding groups without motifs that we do not include in Figure 4. For instance, the concept of group closeness, studied by Chen et al. [22], is later adapted by other researchers to improve the approximation guarantee. Bergamini et al. [10] then scale up the existing approximation, and Angriman et al. [5] offer an algorithm that is one to two orders of magnitude faster, based on a new local search heuristic. Subsequently, Angriman et al. [3] provide a constant approximation for the same local search algorithm.
@@
Nevertheless, there are few exceptions where group centrality metrics are applied to real-world scenarios. Specifically, we observe significant applications in biological networks, particularly in protein-protein interaction networks (refer to [99, 19, 81, 20]). These studies utilize group structures such as star and representative sets in analyzing protein-protein interaction networks to determine essential proteins for the survival of the organisms.
@@ 5 Future work
We observe that the OR community has seen significant advancements on group centrality in the last decade as discussed in Section 3. In addition, we provide additional analysis to provide more depth insights in Section 4. Yet, there are several unexplored areas that might of interest to optimization experts. As recently introduced by Matsypura et al. [60], both walk and path group centrality are evaluated primarily from the largest degree perspective. Thus, there is a research gap for these novel group centrality metrics with respect to closeness and betweenness centrality that we believe needs to be addressed.
Another research direction lies in the domain of tree structures within graphs. Trees, defined as connected graphs where there exists a unique path between each node pair, as a special type of graph, could serve as a basis for developing new group centrality measures. While the concept of spanning tree centrality exists in literature, it differs from our understanding in this paper. The existing studies focus on the probability of an edge being used in a uniformly sampled spanning tree [94, 45]. Identifying “central” trees could serve in application as developing communication protocols; their relationships to connected dominating sets [43, 95] could also be further investigated.
@@
Newman [70] introduces a new measure of betweenness centrality based on random walks. This measure considers the contribution of all paths, not just the shortest ones, while placing more weight on the shortest paths and considering generating random walks (see Noh and Rieger [72]). Later, Gillani et al. [41] extend this idea to group to group version where centrality of a node is measured through a set of source and a set of sink nodes. They indicate that “We look forward to extend random walk based centrality definition to the group centrality setting in which we measure the centrality of groups rather than individual nodes.". We believe that this might be another avenue for further research. Even though this metric might sounds similar to the GED-centrality proposed by Angriman et al. [4], we note that random-walk centrality is more related to algebraic-based metrics including the Laplacian equations.
As for the solution techniques, we note that there is no study proposing exact solution approaches for groups without motifs. Investigating this could lead to significant contributions. Additionally, there is a clear path towards exploring probabilistic group centrality metrics, such as examining the problem of maximum probabilistic clique [63] from a group centrality perspective. Assuming static, unchanging conditions in large-scale network can be a significant limitation in real-world scenarios where uncertainty and variability are inherent. That is why exploring this would be important to understand network dynamics better.
Considering that current models in OR literature are predominantly deterministic, introducing stochastic optimization techniques, like two-stage stochastic models, could be a valuable addition to the field. Such models can help in identifying influential groups that are crucial for the robustness and efficiency of a network under probabilistic conditions. This approach could lead to more resilient and adaptive network designs, thereby enhancing decision-making processes in areas such as infrastructure planning, supply chain and logistics. Additionally, considering the fact that for many networks we do not have a full picture of all the existing nodes and edges (see, e.g., illicit networks which have attracted significant recent interest [52, 6]), stochastic extensions to group centrality can lead to more realistic identification of key actors in networks with hidden elements.
@@
We start from a very important work by Matsypura et al. [60]: identifying paths and walks of high degree can have applications in using autonomous vehicles to monitor a number of targets (by approaching them). As autonomous vehicles are becoming more prominent in applications such as humanitarian relief operations, last mile deliveries, health visits, etc., we believe that the study of closeness metrics from the perspective of walks, paths, and tours/cycles will be an important piece of the puzzle.
Nodal centrality has been prominent in interesting new societal problems, such as investigating criminal and illicit networks [11, 42]. That said, group centrality metrics present a unique opportunity to identify many key illicit actors at once. Moreover, we mentioned earlier the need for more stochastic measures of group centrality. These would be particularly interesting in this context, due to the large number of hidden or obscured entities and relationships.
Additionally, a number of researchers have investigated criticality and centrality from the perspective of disaster management (mostly evacuation). Continuing on that path, we believe it is important to include refueling constraints (see, e.g., Vogiatzis and Kontou [100]) in the calculation of group centrality when applied to evacuation, especially considering the scarcity of gas (and soon other alternative fuels) during an emergency. Finally, another avenue for research is in pedagogy and the formation of groups in the classroom (see, for example, Hill and Peuker [46]). Utilizing group centrality to identify groups with “natural leaders” or groups that satisfy a specific connectivity constraint (e.g., high degrees to other groups or mentors, high closeness to some mentor, etc.) can improve team formation and lead to improved outcomes inside and outside the classroom.